OpenStudy (anonymous):

solve the equation -4(4n+2)=4 the solution is n=??

OpenStudy (anonymous):

-.75

thomaster (thomaster):

-4(4n+2)=4 -16n-8=4 -16n=12 n= 12/-16 = -0.75

thomaster (thomaster):

lol it took 8 minutes to send an answer...